Find the solution of the x in the following x: -5/9 =10/x

Grade:8

2 Answers

Arun
6 years ago
in  – 5/ 9 = 10 /x
if we cross multiply, we get
 – 5*x = 9*10
 – 5x = 90
x = – 90/5
x = –18        answer
hope this helps
Hemakesh Reddy
5 years ago
And is - 18 cross multiply
 
-5/9=10/x
-5 x=90
X=-90/5= -18
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
Bye freind
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
HjhhjkkdgFatsykdufitHddiggjfsullflydktfdfjsoysgjdysdulfoydjgdgxyd